# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v1

@@ II-B Uplink Transmission and Spectral Efficiency
Let $x_{t}$ denote the unit power uplink data signal transmitted by the UE $t$, satisfying $\mathbb{E}{|x_{t}|^{2}}=1$. The uplink normalised signal-to-noise power ratio corresponding to the UE $t$ and additive white Gaussian noise are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively. The uplink signal $\textbf{y}^{u}_{m}\in\mathbb{C}^{A\times 1}$ received at the $m$-th AP is expressed as:
$\textbf{y}^{u}_{m}=\sum_{t\in\mathcal{T}}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
In decentralized data detection, unlike centralized approach where each AP sends received signal and channel estimates to CPU, each AP performs data detection for its served UEs using a local linear received combining vector. Following the proposed decentralized decoding architecture, each AP computes its own interference-suppressing local weight based on the local signal-to-interference-plus-noise ratio (SINR). These weights are applied to the locally detected data symbols before being sent to the CPU. This differs from LSFD, where local data estimates and channel estimates are transmitted to the CPU for centralized processing.
The locally weighted data symbol $\hat{x}_{mt}$ for the UE $t$ at the AP $m$ is given by
@@
where $\mathbf{v}_{m{i_{t}}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for the UE $t$ at the AP $m$, and $a_{mt}$ is the corresponding local weight. These weighted local estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
$\hat{x}_{t}=\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m}.$ (3)
@@
$\hat{x}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}+\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\tilde{\textbf{g}}_{mt}+\sum_{k=1,k\neq t}^{T}\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}x_{k}+\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.$ (4)
Here, the first term is the desired signal for the UE $t$, the second term is interference term due to imperfect channel estimation, the third term is interference term due to all other UEs and last term the noise. Based on the decomposition in (4), the achievable uplink spectral efficiency is obtained using the bounding technique and the Shannon capacity lower bound [17, 18, 16] is given by Theorem 1.
@@
[16, 13] : A lower bound on the uplink ergodic capacity for UE $t$ is given by
@@
where $L_{u}$ is $\Big(\frac{1-\frac{L_{p}}{L_{c}}}{2}\Big)$ and the SINR for the UE $t$ is defined by:
[…]
@@
$\begin{split}&\textsc{DS}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}\right\}\\ &\textsc{BU}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}(\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mt}-\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mt}\right\})\\ &\textsc{UI}_{tk}=\textsc{PC}_{tk}=\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}\\ &\textsc{GN}_{t}=\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m},\end{split}$
where $\mathcal{P}_{i_{t}}$ represents the subset of UEs sharing the pilot $i_{t}$. Also, the SE lower bound, in (5), is valid regardless of the combining scheme used.
@@ III Adaptive Combining and Local Weights Analysis
In this section, we analyze the proposed generalized combining schemes, G-PFZF and G-PWPFZF, from both performance and computational cost perspectives. We first derive closed-form expressions for the uplink SE and local combining weights for both schemes. These expressions provide analytical insights into how the proposed framework improves performance while maintain scalability. Furthermore, we evaluate the computational and fronthauling costs associated with each scheme, highlighting the significant reductions achieved through decentralized decoding system design.
In [13], the combining vectors for any UE $t$ is dependent on $c_{mt}$, which is unique to each UE. As a result, generating a distinct ZF vector for every UE entails a number of multiplications that scales linearly with the total number of UEs. To reduce computational cost, we seek an alternative formulation in which the number of multiplications scales with the number of orthogonal pilot sequences, rather than with the number of UEs.
@@ III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
In the proposed G-PFZF combining scheme, each AP independently designs a combining vector for each of its pilots and pilot-sharing UEs use the same pilot, as APs are not able to distinguish among individual UEs sharing the same pilot. The design of combining vectors depends on whether a pilot is strong or weak. Thus, each AP categorizes the pilots into two groups, strong and weak based on the local sum SE optimization. Unlike traditional PFZF, G-PFZF does not forces the APs to select atleast one pilot as strong and allows the APs to choose the combining schemes among PFZF or MR or FZF.
For the AP $m$, a set of strong pilots is denoted by $\mathcal{S}_{m}$ and a complementary set of weak UEs as $\mathcal{W}_{m}$. Let $L_{\mathcal{S}_{m}}$ denote the number of distinct strong pilots used in the set $\mathcal{S}_{m}$, and define the corresponding set of pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}=\{r_{m,1},r_{m,2},\dots,r_{m,L_{\mathcal{S}_{m}}}\}$. The matrix $\mathbf{\bar{G}}_{m}\in\mathbb{C}^{A\times L_{p}}$ contains the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilot sequences, whereas the strong pilots utilize only $L_{\mathcal{S}_{m}}$ of them. To isolate the relevant columns of $\mathbf{\bar{G}}_{m}$, we define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}=[\mathbf{e}_{r_{m,1}},\dots,\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]\in\mathbb{C}^{L_{p}\times L_{\mathcal{S}_{m}}}$, where $\mathbf{e}_{r}$ denotes the $r$-th column of the identity matrix $\mathbf{I}_{L_{p}}$. The resulting product $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$ yields a full-rank matrix containing only the $L_{\mathcal{S}_{m}}$ linearly independent columns corresponding to the strong pilots. For a given pilot $i\in\mathcal{S}_{m}$, let $j_{mi}\in\{1,2,\dots,L_{\mathcal{S}_{m}}\}$ denote the index of the pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}\in\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}=\mathbf{e}_{i}$.
As discussed earlier, the effective channel matrix corresponding to the strong-pilot UEs is represented as $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$. Following the zero-forcing principle, the pseudo-inverse of this matrix can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$. To obtain the combining vector associated with a specific pilot, this pseudo-inverse term is multiplied by $\bm{\varepsilon}_{j_{mi}}$ and subsequently normalized. Thus, the local ZF combining vector for pilot $i\in\mathcal{S}_{m}$ at AP $m$ is given by:
@@
$\mathbf{v}_{mi}^{\text{MR}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{\mathbb{E}\Big\{\Big\|\mathbf{\bar{G}}_{m}\mathbf{e}_{i}\Big\|^{2}\Big\}}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{A\theta_{mi}}},$ (8)
where $\theta_{mi}=\mathbb{E}\big\{|[\mathbf{\bar{G}}_{m}\mathbf{e}_{i}]_{l}|^{2}\big\}=\big(\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i}\bm{\psi}_{i_{k}}\right|^{2}+1\big)$. Following [19, 20, 21], the normalization term in (7), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix, $\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$, with $A$ degree of freedom satisfying $A-1\geq L_{\mathcal{S}_{m}}$, can be rewritten as
$\sqrt{\mathbb{E}\Big\{\bm{\varepsilon}^{H}_{j_{mi}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\Big\}}=\sqrt{\mathbb{E}\Big\{\big[(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\big]_{j_{mi},j_{mi}}\Big\}}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi}}},$ (9)
Also, the G-PFZF scheme suppresses the interference among the UEs whose pilot lies in $\mathcal{S}_{m}$ by sacrificing the $L_{\mathcal{S}_{m}}$ degree of freedom from the total $A$ degrees to boost the desired signal. Therefore, for any UE $t$ such that $i_{t}\in\mathcal{S}_{m}$
@@
$\begin{split}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},&\text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (11)
@@
$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}_{mk}}\}=\begin{cases}\sqrt{A\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (12)
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\mathbb{E}\Big\{\Big|(\frac{1}{\sqrt{A\theta_{mi_{t}}}c_{mt}}\hat{\mathbf{g}}^{H}_{mt}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (13)
@@
By using the combining vectors in (7) and (8), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by:
@@
where $\delta_{mi_{t}}$ is defined as:
@@
The derivation of SINR expression is given in Appendix -A.
Local Weight Design for G-PFZF Scheme: A fundamental limitation of the conventional LSFD architecture is its reliance on centralized optimization [13]. The optimal LSFD weights are calculated at the CPU using knowledge of global channel statistics, which creates a significant fronthaul signaling overhead and computational burden, thus undermining the scalability of the network. In this work, we break this dependency by introducing a novel decentralized decoding architecture where each AP $m$ independently computes a local interference-suppressing weight $a_{mt}$ for each UE $t$, based solely on its local channel information. Since these weights are derived from the LSFCs, they evolve much slower than the small-scale fading components, ensuring long-term stability and minimal operational overhead.
@@
The main idea behind designing decoding weights is that it should amplify the data signal received from AP with strong local desired signal component and suppress those dominated by interference. In LSFD design, weights are computed through coordination among the APs, specifically the coherent interference terms. In contrast, the proposed decentralized framework eliminates inter-AP coordination, requiring each AP to determine its weight independently. Accordingly, an AP with a higher local SINR is given a larger weight than one with a lower SINR in order to enhance the global SINR. For simplicity, the local weight is chosen to be equal to the local SINR.
The closed-form expression of the local SINR for the UE $t$ at the AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -A. Accordingly, the closed-form of the local SINR or local weight with respect to the AP $m$ of the UE $t$ is given by:
@@
The numerator represents the desired signal strength for the desired UE $t$ at the AP $m$, capturing the gain achieved through the chosen combining strategy. The denominator is designed to penalize the weight by two main sources of impairment: the coherent interference from pilot contamination and the combined power of non-coherent interference and noise.
By applying these weight to its local estimate, each AP independently produces a ”soft” decision that is already pre-optimized for local interference suppression. This design ensures that the signals forwarded to the CPU require no further large-scale fading decoding, thereby completely eliminating the associated fronthaul overhead for channel statistics and the computational cost of centralized weight calculation. The result is a radically simplified and highly scalable uplink processing architecture that retains the performance benefits of distributed APs design.
@@
The foundation of this optimization is the local SINR expression given in (16). As $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$ and substituting $L_{\mathcal{S}m}$ into (16) makes the dependency explicit:
@@
This expression reveals a fundamental trade-off governed by the size of the strong pilot set, $L_{\mathcal{S}m}$. Allocating a pilot to the strong group (i.e., setting $\delta_{mi}=1$) employs ZF to suppress the non-coherent interference from that pilot’s UEs as well as other strong pilots UEs. However, this comes at the cost of a reduced array gain, $(A-L_{\mathcal{S}_{m}})$, for all UEs served by the AP, which diminishes the desired signal power. Conversely, classifying a pilot as weak preserves the full array gain but forgoes any interference suppression for its UEs. This inherent trade-off between maximizing signal strength and minimizing interference makes the grouping problem non-trivial and renders a fixed, network-wide threshold highly suboptimal.
To navigate the above mentioned trade-off, the AP $m$ independently solves a local optimization problem. The objective is to select the binary assignment variables $\delta_{mi}\in\{0,1\}$ for all pilots that maximize the sum of the local SEs for all UEs, as defined by
@@
where $\bm{\delta}_{m}=[\delta_{m1},\delta_{m2},\cdots,\delta_{mL_{p}}]^{\intercal}$. The constraint (18c) represents the necessary condition of Wishart matrix for interference suppression. Also, intuitively, this would otherwise lead to invalid negative values for the desired signal power and coherent interference in the SINR expression.
The optimization problem in (18) is a Mixed-Integer Non-Linear Program (MINLP), a class of problems known for its exponential computational cost. To develop a more tractable formulation, we begin by relaxing the binary constraint (18b) into a continuous one:
$0\leq\delta_{mi}\leq 1\ \forall i.$ (19)
This relaxation transforms the original MINLP into a Non-Linear Program (NLP). To ensure that the solutions of this new NLP satisfy the original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the following complementary constraint for each $i$:
$\delta_{mi}-\delta^{2}_{mi}\leq 0.$ (20)
@@
$\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log(1+\text{SINR}^{\text{G-PFZF}}_{mt}),$ (21a)

### v2

@@ II-B Uplink Transmission and Spectral Efficiency
Let $x_{t}$ denote the unit power uplink data signal transmitted by UE $t$. The uplink normalised data power corresponding to UE $t$ and additive white Gaussian noise are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively. The uplink signal received at the $m$-th AP is:
$\textbf{y}^{u}_{m}=\sum\nolimits_{t=1}^{T}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
In decentralized data detection, unlike centralized approach where each AP sends received signal and channel estimates to CPU, each AP performs data detection for its served UEs using a local linear received combining vector. Following the proposed decentralized decoding architecture, each AP computes its own interference-suppressing local weight based on the local signal-to-interference-plus-noise ratio (SINR). These weights are applied to the locally detected data symbols before being sent to the CPU. This differs from LSFD, where local data and channel estimates are transmitted to the CPU for centralized processing. The locally weighted data symbol $\hat{x}_{mt}$ for UE $t$ at AP $m$ is given by
@@
where $\mathbf{v}_{m{i_{t}}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for UE $t$ at AP $m$, and $a_{mt}$ is the corresponding local weight. These weighted estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
$\hat{x}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m}.$ (3)
@@
$\begin{split}\hat{x}_{t}&=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}+\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\tilde{\textbf{g}}_{mt}\\ &+\sum_{\begin{subarray}{c}k=1\\ k\neq t\end{subarray}}^{T}\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}x_{k}+\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$ (4)
Here, the first term is the desired signal for UE $t$, the second term is interference term due to imperfect channel estimation, the third term is interference term due to all other UEs and last term is the noise. Based on the decomposition in (4), the achievable uplink spectral efficiency is obtained using the bounding technique and the corresponding Shannon capacity lower bound [17, 18, 16] is given by Theorem 1.
@@
[16, 13] : A lower bound on the uplink ergodic SE for UE $t$ is given by
@@
where $L_{u}=\Big(\frac{1-\frac{L_{p}}{L_{c}}}{2}\Big)$ and $\textsc{SINR}_{t}$ is
$\textsc{SINR}_{t}=\frac{p^{u}_{t}|\textsc{DS}_{t}|^{2}}{\!\!\sum\limits_{k=1}^{T}\!\!p^{u}_{k}\mathbb{E}\left\{|\textsc{UI}_{tk}|^{2}\right\}-p^{u}_{t}|\textsc{DS}_{t}|^{2}+\mathbb{E}\left\{|\textsc{GN}_{t}|^{2}\right\}},$ (6)
@@
$\begin{split}&\textsc{DS}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}\right\},\\ &\textsc{UI}_{tk}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk},\\ &\textsc{GN}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$
The SE lower bound, in (5), is valid regardless of the combining scheme used.
@@ III Adaptive Combining and Local Weights Analysis
In this section, we first derive closed-form expressions for the uplink SE and local combining weights for both G-PFZF and G-PWPFZF schemes. We also formulate an optimization problem to partition the pilots into strong and weak.
@@ III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
In [13], the combining vectors for UE $t$ is dependent on $\beta_{mt}$, which is unique to each UE. As a result, generating a distinct combining vector for every UE entails a number of multiplications that scales linearly with the total number of UEs. To reduce computational cost, we propose G-PFZF combining scheme, where each AP independently designs a combining vector for each of its pilots and pilot-sharing UEs use the same pilot, as APs are not able to distinguish among individual UEs sharing the same pilot. The design of combining vectors depends on whether a pilot is strong or weak. Thus, each AP categorizes the pilots into two groups, strong and weak based on the local sum SE optimization. Unlike traditional PFZF, G-PFZF does not force the APs to select atleast one pilot as strong and allows the APs to choose the combining schemes among PFZF or MR or FZF.
Remark 2: In our proposed combining schemes, a strong pilot indicates that a ZF combining vector is assigned to that pilot (and its corresponding UEs), while a weak pilot is assigned an MR combining vector.
For AP $m$, a set of strong pilots is denoted by $\mathcal{S}_{m}$ and a complementary set of weak UEs as $\mathcal{W}_{m}$. Let $L_{\mathcal{S}_{m}}$ denote the number of distinct strong pilots used in set $\mathcal{S}_{m}$, and define the corresponding set of pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}{=}\{r_{m,1},r_{m,2},{\dots},r_{m,L_{\mathcal{S}_{m}}}\}$. The matrix $\mathbf{\bar{G}}_{m}{\in}\mathbb{C}^{A\times L_{p}}$ contains the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilot sequences, whereas the strong pilots utilize only $L_{\mathcal{S}_{m}}$ of them. To isolate the relevant columns of $\mathbf{\bar{G}}_{m}$, we define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}{=}[\mathbf{e}_{r_{m,1}},{\dots},\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]{\in}\mathbb{C}^{L_{p}\times L_{\mathcal{S}_{m}}}$, where $\mathbf{e}_{r}$ denotes the $r$-th column of the identity matrix $\mathbf{I}_{L_{p}}$. The resulting product $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$ yields a full-rank matrix containing only $L_{\mathcal{S}_{m}}$ linearly independent columns corresponding to the strong pilots. For a given pilot $i{\in}\mathcal{S}_{m}$, let $j_{mi}{\in}\{1,2,\dots,L_{\mathcal{S}_{m}}\}$ denote the index of pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}{\in}\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}{=}\mathbf{e}_{i}$.
Following the zero-forcing principle, the pseudo-inverse of this matrix can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$. To obtain the combining vector associated with a specific pilot, this pseudo-inverse term is multiplied by $\bm{\varepsilon}_{j_{mi}}$ and subsequently normalized. Thus, the local ZF combining vector for pilot $i{\in}\mathcal{S}_{m}$ at AP $m$ is given by:
@@
$\mathbf{v}_{mi}^{\text{MR}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{\mathbb{E}\big\{\big\|\mathbf{\bar{G}}_{m}\mathbf{e}_{i}\big\|^{2}\big\}}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{A\theta_{mi}}}.$ (8)
Following [19, 20, 21], the normalization term in (7), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix ($\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$) with degree of freedom satisfying $A\geq L_{\mathcal{S}_{m}}+1$, can be rewritten as
${\mathbb{E}\Big\{\bm{\varepsilon}^{H}_{j_{mi}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\Big\}}=\frac{1}{{(A-L_{\mathcal{S}_{m}})\theta_{mi}}},$ (9)
Also, the G-PFZF scheme suppresses the interference among the UEs whose pilot lies in $\mathcal{S}_{m}$ by sacrificing $L_{\mathcal{S}_{m}}$ degrees of freedom from total $A$ degrees to boost the desired signal. Therefore, for any UE $t$ such that $i_{t}\in\mathcal{S}_{m}$
@@
$\begin{split}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (11)
@@
$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}{\mathbf{\hat{g}}_{mk}}\}=\begin{cases}\sqrt{A\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (12)
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (13)
@@
By using the combining vectors in (7) and (8), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by
@@
where $\mathcal{P}_{i_{t}}$ represents the subset of UEs sharing pilot $i_{t}$ and $\delta_{mi_{t}}$ is defined as:
@@
The derivation of the SINR expression is given in Appendix -A.
Local Weight Design for G-PFZF Scheme: A fundamental limitation of the conventional LSFD architecture is its reliance on centralized optimization [13]. The optimal LSFD weights are calculated at the CPU using knowledge of global channel statistics, which creates a significant fronthaul signaling overhead and computational burden, thus undermining the scalability of the network. In this work, we break this dependency by introducing a novel decentralized decoding architecture where each AP $m$ independently computes a local interference-suppressing weight $a_{mt}$ for each UE $t$, based solely on its local channel information. Since these weights are derived from the LSFCs, they evolve much slowly than the small-scale fading components, ensuring long-term stability and minimal operational overhead.
@@
The main idea behind designing decoding weights is that it should amplify the data signal received from the APs with strong local desired signal component and suppress those dominated by interference. In LSFD design, weights are computed through coordination among the APs, specifically the coherent interference terms. In contrast, the proposed decentralized framework eliminates inter-AP coordination, requiring each AP to determine its weight independently. Accordingly, an AP with a higher local SINR is given a larger weight than the one with a lower SINR in order to enhance the global SINR. For simplicity, the local weight is chosen to be equal to the local SINR.
The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -A. Accordingly, the closed-form of the local SINR or local weight with respect to AP $m$ of UE $t$ is given by:
@@
The numerator represents the desired signal strength for the desired UE $t$ at AP $m$, capturing the gain achieved through the chosen combining strategy. The denominator is designed to penalize the weight by two main sources of impairment: the coherent interference from pilot contamination and the combined power of non-coherent interference due to channel estimation error and noise.
By applying these weight to its local estimate, each AP independently produces a soft decision that is already pre-optimized for local interference suppression. This design ensures that the signals forwarded to the CPU require no further large-scale fading decoding, thereby completely eliminating the associated fronthaul overhead for channel statistics and the computational cost of centralized weight calculation. The result is a radically simplified and highly scalable uplink processing architecture that retains the performance benefits of distributed APs design.
@@
The basis of this optimization is the local SINR expression given in (16). As $L_{\mathcal{S}m}{=}\sum_{i=1}^{L_{p}}\delta_{mi}$ and substituting $L_{\mathcal{S}m}$ into (16) makes the dependency explicit as given in (17).
@@
This expression reveals a fundamental trade-off governed by the size of the strong pilot set, $L_{\mathcal{S}m}$. Allocating a pilot to the strong group ( $\delta_{mi}{=}1$) employs ZF to suppress the non-coherent interference from that pilot’s UEs as well as other strong pilots UEs. However, this comes at the cost of a reduced array gain, $(A-L_{\mathcal{S}_{m}})$, for all UEs served by the AP, which diminishes the desired signal power. Conversely, classifying a pilot as weak preserves the full array gain but forgoes any interference suppression for its UEs. This inherent trade-off between maximizing signal strength and minimizing interference makes the grouping problem non-trivial and renders a fixed, network-wide threshold highly suboptimal.
To navigate the above mentioned trade-off, AP $m$ independently solves a local optimization problem, which is formulated as
@@
where $\bm{\delta}_{m}=[\delta_{m1},\delta_{m2},\cdots,\delta_{mL_{p}}]^{\intercal}$. The constraint (18c) represents the necessary condition of Wishart matrix for interference suppression. Also, intuitively, this would otherwise lead to invalid negative values for the desired signal power and coherent interference in the SINR expression. The optimization problem in (18) is a Mixed-Integer Non-Linear Program (MINLP), a class of problems known for its exponential computational cost. To develop a more tractable formulation, we begin by relaxing the binary constraint (18b) into a continuous one:
$0\leq\delta_{mi}\leq 1,\ \forall i.$ (19)
To ensure that the solutions of this new NLP satisfy the original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the following complementary constraint:
$\delta_{mi}-\delta^{2}_{mi}\leq 0,\ \forall i.$ (20)